OpenStudy (anonymous):

find the absolute maximum of 144(x^3)(y^2)(1-x-y) in the first quadrant of the xy-plane

OpenStudy (anonymous):

I don't understand what you are saying... do I just derive with respect to x, and then plug it into the derivative with respect to y?

OpenStudy (anonymous):

the answer comes to be 1/3 at (1/2, 1/3), but I don't know how to get there..

OpenStudy (turingtest):

we need to get the region... since we are in the first quadrant we have\[144x^3y^2(1-x-y)\ge0\]um... thinking how to graph that :P

OpenStudy (turingtest):

maybe we can just ignore the boundary at the moment and try to find some relative extrema for now...

OpenStudy (anonymous):

How did you get the gradient to equal that? I've been trying to work it out and i have no idea how you got a 4x in there and how y gained an exra power..

OpenStudy (turingtest):

yeah I think that grad is wrong too...

OpenStudy (turingtest):

but I'm eating and am too lazy to work it out lol

OpenStudy (anonymous):

LOL, yeah, i'm trying to work this out, and I'm puzzled..

OpenStudy (anonymous):

\[\left\{-144 x^2 y^2 (4 x+3 y-3),-144 x^3 y (2 x+3 y-2)\right\} \] Sorry, I started with the wrong function.

OpenStudy (anonymous):

Oh alright, but what i'm still now seeing is how you got 4x, or 3y.

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

we need to get the region... since we are in the first quadrant we have 144x3y2(1−x−y)≥0 um... thinking how to graph that :P Try x =2 and y=2

OpenStudy (anonymous):

So the function goes to -Infinity when x and y go to +Infinity

OpenStudy (anonymous):

Before factoring the gradient looks like \[\left\{432 x^2 y^2 (-x-y+1)-144 x^3 y^2,288 x^3 y (-x-y+1)-144 x^3 y^2\right\} \] Factor each component correctly.

OpenStudy (anonymous):

You should find that the absolute maximum of f is \[ \frac 1 3\]

OpenStudy (anonymous):

Here is a graph of your function near where it attains its absoulte Maxinun

OpenStudy (anonymous):

Thank you so much eliassaab! I'm fanning you :)

OpenStudy (anonymous):

You are welcome.